A stationary ball drops from the bleachers and reaches the ground at 11m/s. What is the height of the bleachers? Round to the nearest whole number


1
Expert's answer
2021-02-15T17:38:45-0500

By energy conservation law:

"mgh=mv^2\/2"


"h=\frac{v^2}{2g}=\frac{11^2}{2\cdot9.8}=6m"